# Mining Your Own Secrets: Diffusion Classifier Scores for Continual Personalization of Text-to-Image Diffusion Models

### v2

@@ 3.2 DC scores for Function-space Consolidation
Fig. 3(b) and Algo. 3 outline the working of DSC. Similar to EWC, we compute DC scores using three concepts: the readily available prior concept $\mathbf{c}^{0}$, the $n^{\text{th}}$ task concept $\mathbf{c}^{n}$ learned by the first teacher, and a previous task concept $\mathbf{c}^{j<n}$ learned by the second teacher. The teacher LoRAs should thus generate low-entropy DC scores for their corresponding concepts. The student LoRA is trained to match the DC score distributions from both teachers by minimizing the cross-entropy $H$ for its parameters (Caron et al., 2021). Lastly, while DC scores encode class-specific information for DSC, our primary goal lies in improving the generation quality. We find that relying solely on discriminative DC scores for consolidation is insufficient for noise estimation (see App. fig. 12). Hence, we also incorporate a noise score-matching loss $\mathcal{L}_{\text{MSE}}$ into our DSC learning objective:
@@ 4 Experiments
Implementation. We use the Stable Diffusion v1.4 (Rombach et al., 2022) as our backbone based on the Diffusers library (von Platen et al., 2022). Following CD (Kumari et al., 2023), we train all the models for $1000$ iterations on all but the Celeb-A setup, where we use 2000 training iterations to capture more fine-grained facial attributes (Smith et al., 2024b). In favor of zero-shot generalization, we finetune our hyperparameters only on the six task sequence of Custom Concept. For both EWC and DSC, we set the number of consolidation iterations to ${1/5}^{\text{th}}$ of that of the training iterations number. For computing DC scores, the temperature $\tau$ is set to 1.0 for all but the teacher LoRA in DSC where we set $\tau$ to 0.05. The cardinality of $\mathbf{c}^{k}$ for EWC is set to 5. The DSC loss weights $\gamma$ and $\lambda$ are set to 0.1 and 1.5, respectively. We detail on implementation and hyperparameters in App. H.
@@ 4.1 Results
4.2 Long task sequence
We study the scalability of our proposed methods to a sequence of 50 concepts chosen at random from the Custom Concept dataset (Kumari et al., 2023), with variable number of training images per concept. To avoid any learning bias from large early tasks with more training images, we pick all 50 tasks at random rather than adding them over our six tasks sequence. Fig. 5 compares the results of C-LoRA, LoRA EWC, and Ours (EWC DC). Similar to Smith et al. (2024b), we find that the performance of C-LoRA saturates as the number of tasks grow. Instead, applying EWC on the LoRA parameters emerges as a better performer on the long run. EWC with DC retains the performance particularly on the latter ($>35$) tasks (see App. F.1 for qualitative comparison).
@@ 4.3 Ablation studies
Impact of the number of concepts $k$ for DC scores computation. We study how the number of randomly sampled previous task concepts effects the performance of EWC DC (see App. fig. 17). We notice that excluding the common prior concept $\mathbf{c}^{0}$ in DC scores computation, i.e., $k=2$, leads to the worst performance overall. All setups with $k>2$ perform similar until task 2 as there is only one available previous concept. From task 3 onward, $k=3$ still gets to sample only one previous concept per iteration while $k=7$ can use all previous concepts at each step. We find that the performance of EWC DC saturates as $k$ increases beyond $5$. This is because not all previous concepts carry useful discriminative information for reliable DC scores. We use $k=5$ throughout.
@@ 5 Conclusion
In this paper, we propose continual personalization of pretrained text-to-image diffusion models using their inherent class-conditional density estimates, i.e., Diffusion classifier (DC) scores. Namely, we alleviate forgetting using DC scores as regularizers for parameter-space and function-space consolidation. We design practical considerations for efficiently deriving the DC scores during training. We show the superior performance of our methods through extensive quantitative and qualitative analyses across diverse CL task lengths. Additionally, we show the compatibility of our method for the parameter-efficient VeRA (Kopiczko et al., 2024) and for multi-concept generation (Kumari et al., 2023). We hope that our work paves the general way for leveraging DC scores in personalization of pretrained conditional diffusion models.
@@ Appendix B C-LoRA with sparsity constraint on Task-1 LoRA parameters
Figure 9: C-LoRA with sparsity constraint on the first task: To overcome the catastrophic forgetting of first task in C-LoRA (Smith et al., 2024b) (see Sec. 3.1), we consider restricting LoRA weight updates during first task by incorporating a sparsity (L1 norm) constraint into the training objective for the first task. However, as shown in Fig. 9(a) and 9(b), this merely results in the degenerate solution shifted by one task, i.e., now the task 2 weights (instead of task 1) undergo significant updates, which in turn causes $\mathcal{L}_{\text{forget}}$ for task 3 to decrease throughout training (as most of the task 2 spots get edited). Fig. 9(c) shows the results generated by this model for the first four tasks on our Custom Concept setup. Compared to the results of C-LoRA in Fig. 9(d), now even task 2 image generation (for the pet cat concept) is seen to exhibit catastrophic forgetting.
@@ Appendix D Datasets and their Concepts
For our Google Landmarks v2 (Weyand et al., 2020) setup, we select 10 such geographically diverse waterfall landmarks and download 20 images for each. These landmarks (and their countries) include: Bow falls (Canada), Davis falls (Nepal), Fukuroda falls (Japan), Huka falls (New Zealand), Iguazu falls (Argentina), Korbu falls (Russia), Mang falls (China), Niagara falls (Canada), Rufabgo falls (Russia), and Shin falls (UK).
@@ Appendix F Main Results (Continued)
We study the taskwise performance ($A_{\text{MMD}}$ and KID) evolution of the compared methods on our CL setups of Custom Concept and Landmarks. Fig. 13 and 14 show that overall, our DC score-based variants perform better against their non-DC score-based counterparts as well as against other baselines on every incremental task.
Figure 13: Taskwise performance evolution on Custom Concept CL setup.
Figure 14: Taskwise performance evolution on Landmarks CL setup.
F.1 Generation Results for Long Task Sequence
@@ F.2 Results on Household objects and Real Faces
Figure 17: Effect of the number of concepts $k$ for DC scores computation on the six-tasks Custom Concept setup performance (refer to Algo. 2 on the usage of $k$). Interpretation: $k$ dictates the cardinality of the set $\mathbf{c}^{k}$ that we leverage for preventing the cost of DC scores computation to grow with the number of seen concepts. $\mathbf{c}^{k}$ usually comprises the current $n^{\text{th}}$ task concept $\mathbf{c}^{n}$ and the readily available common prior concept $\mathbf{c}^{0}$. On top of these, we randomly sample $k-2$ previous task concepts that are to be employed in the DC scores computation. As such, all setups with $k>2$ perform similar until task 2 as there is only one available previous concept in the pool to sample from. From task 3 onward, $k=3$ still gets to sample only one previous concept per iteration while $k=5$ can use all previous concepts per iteration (note that our Custom Concept setup has a total of six concepts, and thus the maximum number of previous task concepts is $5$ which is during training on the sixth task). We find that the performance of EWC DC saturates as $k$ increases beyond $5$. This is because not all previous concepts carry useful discriminative information for deriving reliable DC scores. We thus use $k=5$ across all our setups. Lastly, note that DC scores comprise probability distribution over concepts and thus require the minimum number of $2$ concepts for derivation. As a result, we include the pre-trained common prior concept $\mathbf{c}^{0}$ besides the current $n^{\text{th}}$ task concept $\mathbf{c}^{n}$ into $\mathbf{c}^{k}$ on all but the $k=2$ setting. On the latter setting, we sample $1$ previous concept at random per iteration, which is similar to the $k=3$ setup except for the common prior concept included in $\mathbf{c}^{k}$. We find that excluding the common prior concept leads to the worst performance overall. This could be because the prior concept images might help preserve more discriminative semantic information in the DC scores when compared to other downstream task concepts.
@@ G.2 Support for Multi-concept Generation
We study the compatibility of our proposed method for generating multiple custom concepts in the same picture. Figure 18(b) compares the multi-concept results of our method (LoRA EWC DSC DC) with that of C-LoRA (Smith et al., 2024b). Following C-LoRA, we use the prompt style “a photo of $V^{1}$ [$\mathsf{X}$]. Posing with $V^{2}$ [$\mathsf{Y}$]”, where $V^{1}$ and $V^{2}$ are the learnable custom tokens for the concept names $\mathsf{X}$ and $\mathsf{Y}$, respectively. We find that for both the methods, multi-concept generation remains highly sensitive to prompt engineering, and minor prompt changes (replacing ’posing’ with ’together’) can lead to results where one of the concepts is overshadowed by another. We also notice similar effects upon removing the concept names $\mathsf{X}$ and $\mathsf{Y}$ from the prompt. On the other hand, using the concept names in the prompt can lead to interference from the model’s pretrained knowledge, e.g., the waterfall in the background does not always resemble the waterfall from the target images. While both the methods struggle to produce images that preserve the exact style of the target images, we find that our method undergoes relatively lower interference than C-LoRA, which for instance, produces images of the plushy tortoise and sunglasses that differ significantly in color and appearance from the original target images.
@@ H.2.3 Hyperparameters for DSC loss
Appendix I Training Time Complexity Analysis
@@
    For parameter-space consolidation, we rely on online EWC (Schwarz et al., 2018) which maintains a single set of FIM weights that are updated continuously using a running average over tasks. This ensures that our EWC-based framework does not store separate importance weights for each task, and hence, the time complexity scales linearly in the factor of sample size, $\mathcal{\tilde{O}}(|\mathcal{D}|)$. Next, for DC scores computation, we assume a fixed number of conditional forward passes that is proportional to the size of the relevant concept set $\mathbf{c}^{k}$ (see Sec. 3.1). Irrespective of the number of tasks, $\mathbf{c}^{k}$ always stores $m+2$ number of concepts, where $m$ is chosen through grid search and is typically a low number for avoiding confusion from other uninformative classes. Hence, DC scores computation for EWC scales linearly with the number of training samples $\mathcal{\tilde{O}}(|\mathcal{D}|)$. Put together, the time complexity for our parameter-space consolidation framework is: $\mathcal{\tilde{O}}(|\mathcal{D}|)+\mathcal{\tilde{O}}(|\mathcal{D}|)$ = $\mathcal{\tilde{O}}(|\mathcal{D}|)$.
@@
    For function-space consolidation, we rely on a double-distillation framework, which uses two teacher and one student LoRA per consolidation iteration, irrespective of the number of seen tasks. Subsequently, the training time complexity of function-space consolidation remains $\mathcal{\tilde{O}}(|\mathcal{D}|)$. For computing DC scores, we always rely on three conditional forward passes through each of the teachers and the student. As described in Sec. 3.2, these forward passes correspond to the readily available common prior concept $\mathbf{c}^{0}$, and the concepts $\mathbf{c}^{n}$ and $\mathbf{c}^{j<n}$ corresponding to the current task $n$ and the previous task $j<n$ teacher LoRA. Hence, the time complexity of DC scores computation is also $\mathcal{\tilde{O}}(|\mathcal{D}|)$. Overall, the time complexity for our function-space consolidation framework remains: $\mathcal{\tilde{O}}(|\mathcal{D}|)+\mathcal{\tilde{O}}(|\mathcal{D}|)$ = $\mathcal{\tilde{O}}(|\mathcal{D}|)$.
@@
While our method scales better than C-LoRA with the number of tasks (see Table 6), we are nevertheless bounded by the several conditional forward passes needed to derive DC scores for each training minibatch. Despite our proposed considerations for efficient computation of DC scores during consolidation (see Sec. 3.1), the computational overhead for deriving these remains dominant specifically for CL setups with fewer number of tasks. For example, using an RTX A6000, each consolidation iteration for EWC requires $\approx 5.3$s, that for DSC requires $\approx 5.7$s, and that for C-LoRA requires $\approx 0.8$s during finetuning on the task $6$ of our Custom Concept setup. Upon scaling to the $50$ tasks setup, this time gap bridges as the runtime per training iteration of C-LoRA grows to $\approx 3.8$s while that of our methods stay roughly the same ($\approx 5.5$s for EWC, and $\approx 5.72$s for DSC). Therefore, we expect that more clever ways to derive the DC scores during training can effectively reduce the runtime of our methods.

### v3

@@ 3.2 DC scores for Function-space Consolidation
Fig. 3(b) and Algo. 3 outline the working of DSC. Similar to EWC, we compute DC scores using three concepts: the readily available prior concept $\mathbf{c}_{0}$, the $n^{\text{th}}$ task concept $\mathbf{c}_{n}$ learned by the first teacher, and a previous task concept $\mathbf{c}_{j<n}$ learned by the second teacher. The teacher LoRAs should thus generate low-entropy DC scores for their corresponding concepts. The student LoRA is trained to match the DC score distributions from both teachers by minimizing the cross-entropy $H$ for its parameters (Caron et al., 2021). Lastly, while DC scores encode class-specific information for DSC, our primary goal lies in improving the generation quality. We find that relying solely on discriminative DC scores for consolidation is insufficient for noise estimation (see App. fig. 11). Hence, we also incorporate a noise score-matching loss $\mathcal{L}_{\text{MSE}}$ into our DSC learning objective:
@@ 4 Experiments
Implementation. We use the Stable Diffusion v1.4 (Rombach et al., 2022) as our backbone based on the Diffusers library (von Platen et al., 2022). Following CD (Kumari et al., 2023), we train all the models for $1000$ iterations on all but the Celeb-A setup, where we use 2000 training iterations to capture more fine-grained facial attributes (Smith et al., 2024b). In favor of zero-shot generalization, we finetune our hyperparameters only on the six task sequence of Custom Concept. For both EWC and DSC, we set the number of consolidation iterations to ${1/5}^{\text{th}}$ of that of the training iterations number. For computing DC scores, the temperature $\tau$ is set to 1.0 for all but the teacher LoRA in DSC where we set $\tau$ to 0.05. The cardinality of $\mathbf{c}_{k}$ for EWC is set to 5. The DSC loss weights $\gamma$ and $\lambda$ are set to 0.1 and 1.5, respectively. We detail on implementation and hyperparameters in App. H.
@@ 4.1 Results
Long task sequence. We study the scalability of our proposed methods to a sequence of 50 concepts chosen at random from the Custom Concept dataset (Kumari et al., 2023), with variable number of training images per concept. To avoid any learning bias from large early tasks with more training images, we pick all 50 tasks at random rather than adding them over our six tasks sequence. App. fig. 14 compares the results of C-LoRA, LoRA EWC, and Ours (EWC DC). Similar to Smith et al. (2024b), we find that the performance of C-LoRA saturates as the number of tasks grow. Instead, applying EWC on the LoRA parameters emerges as a better performer on the long run. EWC with DC retains the performance particularly on the latter ($>35$) tasks (see App. F.1 for further results).
@@ 4.2 Ablation studies
Impact of the number of concepts $k$ for DC scores computation. We study how the number of randomly sampled previous task concepts effects the performance of EWC DC (see App. fig. 16). We notice that excluding the common prior concept $\mathbf{c}_{0}$ in DC scores computation, i.e., $k=2$, leads to the worst performance overall. All setups with $k>2$ perform similar until task 2 as there is only one available previous concept. From task 3 onward, $k=3$ still gets to sample only one previous concept per iteration while $k=7$ can use all previous concepts at each step. We find that the performance of EWC DC saturates as $k$ increases beyond $5$. This is because not all previous concepts carry useful discriminative information for reliable DC scores. We use $k=5$ throughout.
@@ 5 Conclusion
In this paper, we propose continual personalization of pretrained text-to-image diffusion models using their inherent class-conditional density estimates, i.e., Diffusion classifier (DC) scores. Namely, we alleviate forgetting using DC scores as regularizers for parameter-space and function-space consolidation. We show the superior performance of our methods through extensive quantitative and qualitative analyses across diverse CL task lengths. While DC scores have previously been utilized for test-time classification (Clark & Jaini, 2024), we are the first to advocate that such pre-trained class-specific knowledge of diffusion models can further be reinforced through fine-tuning and can help enhance the performance of downstream generation tasks. We thus hope that our work paves the general way for leveraging DC scores in personalization of pretrained conditional diffusion models.
6 Acknowledgement
We are grateful to Yuhta Takida for their feedback and comments. We would like to acknowledge Sander Dieleman’s blog (Dieleman, 2024) for insights on diffusion distillation.
@@ Appendix A Personalization in Text-to-Image Diffusion Models
How do we obtain and leverage the new word vector $V_{n}^{*}$ in the training process? Following TI (Gal et al., 2023) and CD (Kumari et al., 2023), to personalize our text-to-image diffusion model on a new concept $\mathbf{c}_{n}$, we introduce a new token representing this concept and learn its corresponding word vector $V_{n}^{*}$ by optimizing only this embedding and the $n^{\text{th}}$ task LoRA while keeping the rest of the model’s parameters frozen. To do so, we create prompts for the $n^{\text{th}}$ concept that include the new token (e.g.,“a photo of [$V_{n}^{*}$]”). By inputting these prompts into the model and comparing the generated images with the training images, our standard denoising score matching objective (Eq. 1) measures how well the model reproduces the concept. Minimizing this loss adjusts the word vector $V_{n}^{*}$ so that the model associates the new token with the visual characteristics of $\mathbf{c}_{n}$, enabling it to generate images of the concept when the token is used in prompts. Lastly, as mentioned in Sec. 3, $V_{n}^{*}$ is acquired during the training stage and remains frozen thereafter, i.e., while we perform consolidation using the DC scores.
@@ Appendix B C-LoRA with sparsity constraint on Task-1 LoRA parameters
Figure 8: C-LoRA with sparsity constraint on the first task: To overcome the catastrophic forgetting of first task in C-LoRA (Smith et al., 2024b) (see Sec. 3.1), we consider restricting LoRA weight updates during first task by incorporating a sparsity (L1 norm) constraint into the training objective for the first task. However, as shown in Fig. 8(a) and 8(b), this merely results in the degenerate solution shifted by one task, i.e., now the task 2 weights (instead of task 1) undergo significant updates, which in turn causes $\mathcal{L}_{\text{forget}}$ for task 3 to decrease throughout training (as most of the task 2 spots get edited). Fig. 8(c) shows the results generated by this model for the first four tasks on our Custom Concept setup. Compared to the results of C-LoRA in Fig. 8(d), now even task 2 image generation (for the pet cat concept) is seen to exhibit catastrophic forgetting.
@@ Appendix D Datasets and their Concepts
For our Google Landmarks v2 (Weyand et al., 2020) setup, we select 10 such geographically diverse waterfall landmarks and download 20 images for each. These landmarks (and their countries) include: Bow falls (Canada), Davis falls (Nepal), Fukuroda falls (Japan), Huka falls (New Zealand), Iguazu falls (Argentina), Korbu falls (Russia), Mang falls (China), Niagara falls (Canada), Rufabgo falls (Russia), and Shin falls (the UK).
@@ Appendix F Main Results (Continued)
We study the taskwise performance ($A_{\text{MMD}}$ and KID) evolution of the compared methods on our CL setups of Custom Concept and Landmarks. Fig. 12(a) and 12(b) show that overall, our DC score-based variants perform better against their non-DC score-based counterparts as well as against other baselines on every incremental task.
(a) Taskwise performance evolution on Custom Concept CL setup.
(b) Taskwise performance evolution on Landmarks CL setup.
Figure 12: Taskwise performances.
F.1 Results for Long Task Sequence
@@ F.2 Results on Household objects and Real Faces
Figure 16: Effect of the number of concepts $k$ for DC scores computation on the six-tasks Custom Concept setup performance (refer to Algo. 2 on the usage of $k$). Interpretation: $k$ dictates the cardinality of the set $\mathbf{c}_{k}$ that we leverage for preventing the cost of DC scores computation to grow with the number of seen concepts. $\mathbf{c}_{k}$ usually comprises the current $n^{\text{th}}$ task concept $\mathbf{c}_{n}$ and the readily available common prior concept $\mathbf{c}_{0}$. On top of these, we randomly sample $k-2$ previous task concepts that are to be employed in the DC scores computation. As such, all setups with $k>2$ perform similar until task 2 as there is only one available previous concept in the pool to sample from. From task 3 onward, $k=3$ still gets to sample only one previous concept per iteration while $k=5$ can use all previous concepts per iteration (note that our Custom Concept setup has a total of six concepts, and thus the maximum number of previous task concepts is $5$ which is during training on the sixth task). We find that the performance of EWC DC saturates as $k$ increases beyond $5$. This is because not all previous concepts carry useful discriminative information for deriving reliable DC scores. We thus use $k=5$ across all our setups. Lastly, note that DC scores comprise probability distribution over concepts and thus require the minimum number of $2$ concepts for derivation. As a result, we include the pre-trained common prior concept $\mathbf{c}_{0}$ besides the current $n^{\text{th}}$ task concept $\mathbf{c}_{n}$ into $\mathbf{c}_{k}$ on all but the $k=2$ setting. On the latter setting, we sample $1$ previous concept at random per iteration, which is similar to the $k=3$ setup except for the common prior concept included in $\mathbf{c}_{k}$. We find that excluding the common prior concept leads to the worst performance overall. This could be because the prior concept images might help preserve more discriminative semantic information in the DC scores when compared to other downstream task concepts.
@@ G.2 Support for Multi-concept Generation
We study the compatibility of our proposed method for generating multiple custom concepts in the same picture. Figure 17(b) compares the multi-concept results of our method (LoRA EWC DSC DC) with that of C-LoRA (Smith et al., 2024b) and CD (Kumari et al., 2023). Following C-LoRA, we use the prompt style “a photo of $V^{1}$ [$\mathsf{X}$]. Posing with $V^{2}$ [$\mathsf{Y}$]”, where $V^{1}$ and $V^{2}$ are the learnable custom tokens for the concept names $\mathsf{X}$ and $\mathsf{Y}$, respectively. We find that for both the methods, multi-concept generation remains highly sensitive to prompt engineering, and minor prompt changes (replacing ’posing’ with ’together’) can lead to results where one of the concepts is overshadowed by another. We also notice similar effects upon removing the concept names $\mathsf{X}$ and $\mathsf{Y}$ from the prompt. On the other hand, using the concept names in the prompt can lead to interference from the model’s pretrained knowledge, e.g., the waterfall in the background does not always resemble the waterfall from the target images. While both the methods struggle to produce images that preserve the exact style of the target images, we find that our method undergoes relatively lower interference than C-LoRA, which for instance, produces images of the plushy tortoise and sunglasses that differ significantly in color and appearance from the original target images.
Following our standard evaluation setup (Sec. 4), to quantify the performance for multi-concept generation, we generate 400 images each for the prompts used in Fig. 17(b): “A photo of $V^{1}$ plushie tortoise. Posing in front of $V^{2}$ waterfall” and “A photo of $V^{1}$ plushie tortoise. Posing with $V^{2}$ sunglasses”. In the absence of ground-truth images that incorporate multiple concepts, we quantify the scores by first relying on the ground truth images of the individual concepts in each of these prompts and then averaging out the two scores. Table 6 reports the scores averaged over the two multi-concept generation prompts. We note that these scores are in line with Fig. 17(b) where our EWC DSC DC variant performs significantly better than C-LoRA on all three performance quantification metrics.
Table 6: Results for multi-concept generation on Custom Concept setup
Method KID ($\downarrow$) (x $10^{5}$) $A_{\text{MMD}}$($\downarrow$) (x $10^{3}$) CLIP I2I ($\uparrow$) (x100)
CD 231.7 194.11 48.50
C-LoRA 193.61 130.80 52.45
Ours 163.85 113.42 64.53
@@
(a) 100 iterations
(b) 200 iterations (standard setup)
(c) 300 iterations
(d) 400 iterations
Figure 19: Effect of varying consolidation iterations on FIM estimation: Top-5 Eigenvalue comparison for the FIM estimated with and without DC scores using the learned LoRA parameters for a randomly chosen downsampling layer block of the U-Net. Upon increasing the number of consolidation iterations to 200, we observe larger magnitudes of Eigenvalues, thus indicating that the learned parameters become more certain about the directions of high loss changes from data, i.e., reduced uncertainty in the FIM estimation. However, increasing the number of consolidation iterations beyond 200 leads to a saturation in the data-informed uncertainty estimation for the FIM.
G.3 Compatibility with Stable Diffusion v2.0
Table 7: Custom Concept results for Stable Diffusion v2.0 (avg. over 3 seeds)
C-LoRA 0.05 100.29 158.9 114.5 -103.88 68.10 0.015
EWC 0.05 100.29 141.3 100.97 -96.50 75.81 0.011
Ours: EWC DC 0.05 100.29 134.97 93.59 -88.30 77.29 0.0009
Ours: DSC 0.05 100.29 177.92 155.06 -102.55 75.01 0.003
Ours: DSC EWC 0.05 100.29 139.76 93.71 -91.02 77.90 0.006
Ours: DSC EWC DC 0.05 100.29 126.43 88.54 -82.15 78.22 0.001
Table 8: Landmarks results for Stable Diffusion v2.0 (avg. over 3 seeds)
C-LoRA 0.05 100.32 101.33 62.40 -12.99 80.50 0.009
EWC 0.05 100.32 59.21 38.66 -39.05 85.37 0.002
Ours: EWC DC 0.05 100.32 52.80 32.65 -5.19 87.22 0.001
Ours: DSC 0.05 100.32 116.83 67.32 -65.70 80.01 0.009
Ours: DSC EWC 0.05 100.32 59.40 45.27 -5.71 86.88 0.0007
Ours: DSC EWC DC 0.05 100.32 46.15 29.13 -2.90 88.14 0.084
@@ H.2.1 Number of consolidation iterations
(a) 2000 iterations
(b) 1000 iterations
(c) 400 iterations
(d) 200 iterations
Figure 22: Distribution of uniformly sampled timesteps over a varying total number of training iterations: (a) 2000 finetuning iterations as used by Smith et al. (2024b) for Celeb-A faces, (b) 1000 finetuning iterations as used by Custom Diffusion Kumari et al. (2023) that we adopt for all but the Celeb-A setup, (c) 400 consolidation iterations that we adopt for the Celeb-A setup, (d) 200 consolidation iterations that we adopt for all but the Celeb-A setup. Note that the values for (c) and (d) are chosen per the hyperparameter tuning we perform in App. H.2.1.
@@ H.2.3 Hyperparameters for DSC loss
H.2.4 Hyperparameter for EWC+DSC loss interactions
Figure 27: Effect of $\delta$ and $\gamma$ on performance of DSC EWC DC
Appendix I Training Time Complexity Derivation and Analyses
@@
    For parameter-space consolidation, we rely on online EWC (Schwarz et al., 2018) which maintains a single set of FIM weights that are updated continuously using a running average over tasks. This ensures that our EWC-based framework does not store separate importance weights for each task, and hence, the time complexity scales linearly in the factor of sample size, $\mathcal{\tilde{O}}(|\mathcal{D}|)$. Next, for DC scores computation, we assume a fixed number of conditional forward passes that is proportional to the size of the relevant concept set $\mathbf{c}_{k}$ (see Sec. 3.1). Irrespective of the number of tasks, $\mathbf{c}_{k}$ always stores $m+2$ number of concepts, where $m$ is chosen through grid search and is typically a low number for avoiding confusion from other uninformative classes. Hence, DC scores computation for EWC scales linearly with the number of training samples $\mathcal{\tilde{O}}(|\mathcal{D}|)$. Put together, the time complexity for our parameter-space consolidation framework is: $\mathcal{\tilde{O}}(|\mathcal{D}|)+\mathcal{\tilde{O}}(|\mathcal{D}|)$ = $\mathcal{\tilde{O}}(|\mathcal{D}|)$.
@@
    For function-space consolidation, we rely on a double-distillation framework, which uses two teacher and one student LoRA per consolidation iteration, irrespective of the number of seen tasks. Subsequently, the training time complexity of function-space consolidation remains $\mathcal{\tilde{O}}(|\mathcal{D}|)$. For computing DC scores, we always rely on three conditional forward passes through each of the teachers and the student. As described in Sec. 3.2, these forward passes correspond to the readily available common prior concept $\mathbf{c}_{0}$, and the concepts $\mathbf{c}_{n}$ and $\mathbf{c}_{j<n}$ corresponding to the current task $n$ and the previous task $j<n$ teacher LoRA. Hence, the time complexity of DC scores computation is also $\mathcal{\tilde{O}}(|\mathcal{D}|)$. Overall, the time complexity for our function-space consolidation framework remains: $\mathcal{\tilde{O}}(|\mathcal{D}|)+\mathcal{\tilde{O}}(|\mathcal{D}|)$ = $\mathcal{\tilde{O}}(|\mathcal{D}|)$.
@@
While our method scales better than C-LoRA (Smith et al., 2024b) with the number of tasks (see Table 9), we are nevertheless bounded by the several conditional forward passes needed (depending on the value of $k$) to derive DC scores for each training minibatch. Despite our proposed considerations for efficient computation of DC scores during consolidation (see Sec. 3.1), the computational overhead for deriving these remains dominant specifically for CL setups with fewer number of tasks. For example, using an RTX A6000, each consolidation iteration for EWC requires $\approx 5.3$s, that for DSC requires $\approx 5.7$s, and that for C-LoRA requires $\approx 0.8$s during finetuning on the task $6$ of our Custom Concept setup. As shown in Table 10, scaling to the $50$ tasks setup, this time gap bridges as the runtime per training iteration of C-LoRA grows to $\approx 3.8$s while that of our methods stay roughly the same ($\approx 5.5$s for EWC, and $\approx 5.72$s for DSC for $k=5$). It is worth noting that on the $50^{\text{th}}$ training task, for lower values of $k$, the runtime per training iteration for our methods remain comparable (for $k=3$) or significantly lower (for $k=2$) than that of C-LoRA. Therefore, we expect that more clever ways to derive the DC scores during training can effectively reduce the runtime of our methods.
Table 10: Comparison of training time per iteration (wall clock time in seconds) with varying $k$
Method k = 2 k = 3 k = 5
On $6^{\text{th}}$ task
C-LoRA 0.8
Ours: EWC DC 1.27 3.4 5.3
Ours: DSC DC 1.65 3.61 5.7
On $50^{\text{th}}$ task
C-LoRA 3.8
Ours: EWC DC 1.36 3.57 5.5
Ours: DSC DC 1.8 3.66 5.72